Peacock Corporation has a target capital structure of 70 percent common stock, 5 percent preferred stock, and 25 percent debt. Its cost of equity is 11 percent, the cost of preferred stock is 5 percent, and the cost of debt is 7 percent. The relevant tax rate is 35 percent.

a.

What is the company’s WACC? (Do not round intermediate calculations. Round the final answer to 2 decimal places.)

  WACC %  

   

b.

What is the after-tax cost of debt? (Do not round intermediate calculations. Round the final answer to 2 decimal places.)
Answer #1

Solution:

The formula for calculating the weighted average cost of capital is =

WACC = [ Ke * We ] + [ Kp * Wp ] + [ ( Kd * ( 1- t ) ) * Wd ]

Ke = Cost of equity or common stock ; We = Weight of equity or common stock;

Kp = Cost of preferred stock ; Wp = Weight of preferred stock ;                            

Kd = Cost of debt  ; t = Income tax rate ; Wd = Weight of debt

As per the information available in the question we have

Ke = 11 %    ; We = 70 % = 0.70 ;   Kp = 5 %   ; Wp = 5 % = 0.05 ;         t = 35 % = 0.35   ; Kd = 7 %    ; Wd = 25 % = 0.25

Applying the above values in the formula we have

= [ 11 * 0.70 ] + [ 5 *0.05 ] + [ 7 * ( 1 – 0.35 ) * 0.25 ]

= 7.7 + 0.25 + ( 4.55 * 0.25 )

= 7.7 + 0.25 + 1.1375

= 9.0875

Thus the company’s WACC is = 9.0875 %

= 9.09 % ( when rounded off to two decimal places )

Calculation of after tax cost of debt:

After tax cost of debt = Kd * ( 1- t )

As per the information given in the question we have

Kd = 7 %   ;        t = 35 % = 0.35

= 7 * ( 1 -0.35 ) = 7 * 0.65 = 4.55 %

Thus the after tax cost of debt = 4.55 %

Thus solution to question a = Company's WACC = 9.09 %

Thus solution to question b = After tax cost of debt = 4.55 %